Carlisle Company has been cited and must invest in equipment to reduce stack emissions or face EPA fines of $18,500 per year. An emission reduction filter will cost $75,000 and have an expected life of 5 years. Carlisle's MARR is 10%/yr.

a.  

What is the internal rate of return of this investment?

b.  

What is the decision rule for judging the attractiveness of investments based on internal rate of return?

c.  

Is the filter economically justified?
Answer #1

Initial cost of emission reduction filter = $75,000

Savings in EPA fines = $18,500 per year

MARR = 10%

Life = 5 years

a. What is internal rate of return?

Calculating IRR using Trial & Error Method

Let the rate of interest is 7%. Calculate the PW of the investment at 7%.

PW = -75,000 + 18,500 (P/A, 7%, 5)

PW = -75,000 + 18,500 (4.10020) = 853.7

The PW is positive. Increase the rate of interest to get negative PW. Increase the rate of interest to 8% and calculate the Present Worth.

PW = -75,000 + 18,500 (P/A, 8%, 5)

PW = -75,000 + 18,500 (3.99271) = -1,134.87

Using interpolation,

IRR = 7% + [853.7 – 0 ÷ 853.7 – (-1,134.87)] * 1

IRR = 7.42%

The internal rate of return is 7.42%.

b. Decision Rule

If the IRR > MARR – Select the Investment

If the IRR < MARR – Reject the Investment

c. Is the investment is economically justified?

Answer – No, as the IRR < MARR – Reject the Investment. The investment is not justified.